# Real-time Tracking System with partially coupled sources

### v1

@@ I Introduction
Works [3, 4, 6, 7] investigate status update systems with correlated sources, focusing on age of information (AoI), while works [8, 9] investigate estimation error in such systems. The work [8] derived closed-form expressions of average AoI and state estimation error in a multi-sensor system with correlated sources under an M/M/1/1 queue model. The work [9] derived upper and lower bounds on estimation error using weighted sum AoI and proposed near-optimal AoI-based scheduling for correlated sources. Unlike these works, we consider a pull-based status update system where the scheduler (sink) only has partial knowledge about the source states. In contrast to [8, 9], we characterize the correlation through a coupling factor, which allows for a structured analysis of its impact. Moreover, we solve an optimization control problem to minimize estimation error and transmission cost jointly, whereas [8] characterized a closed-form expression of the average estimation error and [9] does not consider transmission cost.
@@ III-B Belief-MDP Characterization
The state of the sources X is only partially observable at the controller (sink) for decision-making. To address this insufficiency, we utilize the notion of belief [11, Chapter 7], and we cast the above-defined POMDP as a belief-MDP. To this end, we define a belief about the source states that retain the Markov property and capture all information necessary for decision-making.
@@
To solve the belief-MDP problem, we implement the relative value iteration algorithm (RVIA) [^4] For RVIA convergence, the MDP must be unichain and aperiodic [12, Proposition 4.3.2], which is nontrivial to verify. However, numerical results indicate RVI convergence. [12, Proposition 4.3.2]. The RVIA requires a finite belief-state and action space. However, due to the belief element, the belief-state space is infinite. To truncate it, we define a finite space comprising belief trajectories over $N$ transition steps for all action combinations. Beliefs at step $N+1$ are projected onto the closest belief within this space, measured by mean-square error. The belief-state transition probabilities are then updated to account for transitions from step $N$ to the projected beliefs at step $N+1$. The derived policy is referred to as POMDP-based policy.

### v2

@@ I Introduction
Works [3, 4, 6, 7] investigate status update systems with correlated sources, focusing on age of information (AoI), while works [8, 9, 10] investigate estimation error in such systems. The work [8] derived closed-form expressions of average AoI and state estimation error in a multi-sensor system with correlated sources under an M/M/1/1 queue model. The work [9] derived upper and lower bounds on estimation error using weighted sum AoI and proposed near-optimal AoI-based scheduling for correlated sources. Unlike these works, we consider a pull-based status update system where the scheduler (sink) only has partial knowledge about the source states. In contrast to [8, 9], we characterize the correlation through a coupling factor, which allows for a structured analysis of its impact. The work in [10] studied a pull-based status update system with two correlated sources. Unlike our model, where the destination has only partial knowledge about the other sources, the correlation in [10] is captured through probabilistic observation. Specifically, the destination may decode the status of one source upon receiving an update from the other. Due to this fundamental difference, the policies derived in [10] are not applicable to our system. Moreover, we solve an optimization control problem to minimize estimation error and transmission cost jointly, whereas [8] characterized a closed-form expression of the average estimation error and [9] does not consider transmission cost.
@@ III-B Belief-MDP Characterization
The state of the sources X is only partially observable at the controller (sink) for decision-making. To address this insufficiency, we utilize the notion of belief [12, Chapter 7], and we cast the above-defined POMDP as a belief-MDP. To this end, we define a belief about the source states that retain the Markov property and capture all information necessary for decision-making.
@@
To solve the belief-MDP problem, we implement the relative value iteration algorithm (RVIA) [^4] For RVIA convergence, the MDP must be unichain and aperiodic [13, Proposition 4.3.2], which is nontrivial to verify. However, numerical results indicate RVI convergence. [13, Proposition 4.3.2]. The RVIA requires a finite belief-state and action space. However, due to the belief element, the belief-state space is infinite. To truncate it, we define a finite space comprising belief trajectories over $N$ transition steps for all action combinations. Beliefs at step $N+1$ are projected onto the closest belief within this space, measured by mean-square error. The belief-state transition probabilities are then updated to account for transitions from step $N$ to the projected beliefs at step $N+1$. The derived policy is referred to as POMDP-based policy.